Suppose that you hit a tennis ball with an initial speed of 15.4 m/s, at an angle of 48.7
Answer #1

h=vo*t*sin(48.7)-0.5gt^2

2.1 = 11.57t - 4.9t^2

t1 = 2.16

t2 = 0.198

so

d=vo*cos(48.7)*2.16 = 21.95

21.95 - 10 = 11.95

2.16- 0.25 = 1.91

v = d/t = 11.95/1,91 = 6.26 m/s